Calculate the elevation of boiling point when 21.2 g of Potassium Nitride is added to 135mL of water? The boiling point constant of water is 0.52 oC/m
1
Expert's answer
2018-01-23T04:30:25-0500
∆T boil = E·Cm

M(K3N) = 39*3 + 14 = 131 g/mol

n(K3N) = 21.2 g / 131 g/mol = 0.162 mol

Cm(K3N) = 0.162 mol / 0.135 kg = 1.2 mol/kg

∆T boil = 0.52 °C·kg/mol * 1.2 mol/kg = 0.624 °C